1) A particular 12 V car battery can send a total charge of 84 A h (ampere-hours) through a circuit, from one terminal to the other. (a) How many coulombs of charge does this represent? (Hint: current = ) (b) If this entire charge undergoes a change in electric potential of 12 V, how much energy is involved?
Answer #1
Given that the potential of car battery is V = 12 V
The total charge is 84 Ah
----------------------------------------------------------
Now the charge q related to current I as
             I = q /t
             q = I x t
                = 84 A.h
                = 84*60*60s A.s
                = 302400 C  
Since one hour = 60min = 60*60s = 3600s
--------------------------------------------------------
The energy involved
        U = q *V
            = 302400 C *12 V
            = 3628800 J